On January 2,2009,abc company purchased a patent for $38000.The patent has an estimated useful life of 25 years and a20 years legal life.

Required: what entry would the company make at december 31,2009 to recored amortization expense on the patent.


1
Expert's answer
2021-03-30T19:56:17-0400
"\frac {38000}{25}=1520"

"\frac {1520}{365}\times 9=37.48"